OpenStudy (anonymous):

Which of the following is the formula for the geometric sequence? 1.2, 2.4, 4.8, 9.6,... an = 9.6(1/2)n-1 an = 1.2(2)n-1 an = 9.6(2)n-1

OpenStudy (anonymous):

Here is one more: Which of the following is the formula for the geometric sequence? 900, 300, 100, 33 1/3,... an = 900(3)n-1 an = 33 1/3 (3)n-1 an = 900(1/3)n-1

OpenStudy (anonymous):

\[a _{n}=1,2(2)^{n-1}\] \[a _{n}=900(1/3)^{n-1}\]

OpenStudy (anonymous):

\[a _{n}=a(r)^{n-1}\] is the general term where a is the first term